OpenStudy (anonymous):

find the slope intercept equation of the line that has the given characteristics. Slope -7 and y-intercept (0,6)

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

So here is how proper slope intercept is set k? And bear with me I am not always clear so please let me know k?

OpenStudy (anonymous):

so slope intercept is Y equals Mx +B

OpenStudy (anonymous):

So your Y will be 6 for this one as it is the set of cords given.

OpenStudy (anonymous):

y = mx + b where m is your slope and b is your y int

OpenStudy (anonymous):

slope turns into -7 over 1

OpenStudy (anonymous):

Ok? Good so far sweetpea?

OpenStudy (anonymous):

so 6 = -7/1 and you want b

OpenStudy (anonymous):

Which is 0

OpenStudy (anonymous):

k?

OpenStudy (anonymous):

Let me double check my math but that should be right.

OpenStudy (anonymous):

Should be right. Math is right :)

OpenStudy (anonymous):

Let me know if that is right or wrong. k?

OpenStudy (anonymous):

so, it'll be 6=-7x+0? but for my paper it'll have to be y=-7x+0

OpenStudy (anonymous):

Hmmmm might be asking for a different way of doing it. Give me a few to work it out alright?

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

b is 6, not 0. -7 is your m, 6 is your constant. y = -7x + 6

OpenStudy (anonymous):

One of the reasons I detest math. One answer and a thousand bloody ways to do it and each is wrong to someone else's view XD

OpenStudy (anonymous):

6 is your y pete. Almost got it.

OpenStudy (anonymous):

I'm in college and it is actually one of my better areas.

OpenStudy (anonymous):

Here we go sweetpea. As I was saying I had it right :) Write it out for you to see how it goes alright?

OpenStudy (anonymous):

now i'm confused again. lol....so it's not going to be y=-7x+6?

OpenStudy (anonymous):

Hang on sweetpea and I will explain. :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Basic slope intercept form is Y=mx +b What you seek is B

OpenStudy (anonymous):

Slope is -7 and your points are 0.6

OpenStudy (anonymous):

ok, so as of right now we have y=-7x+b still, correct???

OpenStudy (anonymous):

In order to gain b you must input the numbers like this y= 6 mx= 0 and -7

OpenStudy (anonymous):

So you write it out to 6=(0)(-7)+b

OpenStudy (anonymous):

Anything multiplied by zero is zero.

OpenStudy (anonymous):

lol, then let's say i'd hate to go to your college.

OpenStudy (anonymous):

Pete please hold on your answer so we do not confuse her.

OpenStudy (anonymous):

So that leaves us with 6=0+b which point you subtract 0.

OpenStudy (anonymous):

Leaving you with 6=b

OpenStudy (anonymous):

Want to check me? Look up slope intercept form but your answer is 6 sweetpea :)

OpenStudy (anonymous):

may i ask what college you go to?

OpenStudy (anonymous):

Sorry I do not give out that info but it is a good college and again doubt me look it up.

OpenStudy (anonymous):

I aced this section thank you and even then I always double check another source. Did this help you sweetpea?

OpenStudy (anonymous):

yes, once i answer 2 more questions, i will be able to submit my test and will know for sure if it's right or wrong.

OpenStudy (anonymous):

Alright. :)

OpenStudy (anonymous):

i appreciate the help from both of you. :)

OpenStudy (anonymous):

Np hun I always try to help others :)

OpenStudy (anonymous):

i'd like to post my last 2 questions for help so that i can hurry and be done. i'm going to close this one out so that it's not just sitting around.

OpenStudy (anonymous):

Alright.

OpenStudy (anonymous):

lol just saying, it can't be all that great if you're struggling with 8th grade arithmetic. it seems like you can't distinguish between y (your function) and y-int (6).